Insomnia Cookies in Newark is considering a fall promotion that will cost $2500. They expect incremental cash flows to be: Year 1 $600, Year 2 $1300, Year 3 $800, Year 4 $$600. How long will it take them to recoup the initial outlay?
Answer #1

Investment = $2500

Till year 2 cahflows = 600+1300=1900$

Still we need 2500-1900=600 to recoup investment, which will take 600/800=0.75 years of the 3rd year = 2+0.75=2.75 years